OpenStudy (unklerhaukus):

\[(2t^3y-y^3)\text dt-2t^4\text dy=0\]

OpenStudy (unklerhaukus):

\[\frac{\partial M}{\partial y}=2t^3-3y^2\qquad\qquad\frac{\partial N}{\partial t}=-8t^3\] \[\frac{\partial M}{\partial y}\neq\frac{\partial N}{\partial t}\]

OpenStudy (unklerhaukus):

\[R=R(t,y)=t^my^n\] \[\frac{\partial MR(t,y)}{\partial y}=\frac{\partial NR(t,y)}{\partial t}\] \[R\frac{\partial M}{\partial y}+R_yM=R\frac{\partial N}{\partial t}+R_tN\]

OpenStudy (unklerhaukus):

\[t^my^n(2t^3)+nt^my^{n-1}(2t^3y)=t^my^n(-8t^3)+mt^{m-1}y^n(-2t^4)\] \[2t^{m+3}y^n+2nt^{m+3}y^{n}=-8t^{m+3}y^n-2mt^{m+3}y^n\] \[(10+2n+2m)t^{m+3}y^n=0\]

OpenStudy (anonymous):

Can you tell me what topic is this ??

OpenStudy (unklerhaukus):

first order differential equations - exact equations - exact equations with integrating factor of both variables

OpenStudy (unklerhaukus):

i think i've worked it out now

OpenStudy (unklerhaukus):

its a differnt way then i was expecting

OpenStudy (anonymous):

But you are getting the same result.. For this topic I have to google and then I can understand what part Mr Uncle Rocks is doing.. I am unfamiliar with it..

OpenStudy (unklerhaukus):

i wasent getting any result the first time i was getting stuck..

OpenStudy (anonymous):

Suppose I want to search this topic then can you tell me what should I write their so that I can find how to solve these types of problems? Any particular name to search this topic ..

OpenStudy (unklerhaukus):

\[\frac{\partial MR(t,y)}{\partial y}=\frac{\partial NR(t,y)}{\partial t}\] \[\frac{\partial }{\partial y}\left((2t^3y-y^3)t^my^n\right)=\frac{\partial }{\partial t}\left((-2t^4)t^my^n\right)\]\[\frac{\partial }{\partial y}\left(2t^{m+3}y^{n+1}-t^my^{n+3}\right)=\frac{\partial }{\partial t}\left(-2t^{m+4}y^n\right)\]\[2(n+1)t^{m+3}y^{n}-(n+3)t^my^{n+2}=-2(m+4)t^{m+3}y^n\] \[2(n+1)=-2(m+4);\qquad\qquad (n+3)=0\]\[n+1=-m-4;\qquad\qquad n=-3\]\[-2=-m-4;\qquad\qquad m=-2\] \[R(t,y)=t^{-3}y^{-2}\]

OpenStudy (unklerhaukus):

applying the product rule was a mistake,

OpenStudy (anonymous):

Oh that is okay.. I am still understanding it.. take me time to see how you solved it..

OpenStudy (unklerhaukus):

i havent solved it yet, but i have found the indicies of the integrating factor

OpenStudy (anonymous):

Now what to do next??

OpenStudy (unklerhaukus):

\[f(t,y)=\int R(t,y)M\text dt\]

OpenStudy (unklerhaukus):

\[f(t,y)=\int t^{-2}y^{-3}(2t^3y-y^3)\text dt\]\[=\int \left(2ty^{-2}-t^{-2}\right)\text dt\]\[=\frac{t^2}{y^{2}}+\frac{1}{2t}+g(y)\]

OpenStudy (unklerhaukus):

\[f(t,y)=\int R(t,y)N\text dy\]

OpenStudy (anonymous):

Let me try this..

OpenStudy (anonymous):

Don't you think you are wrong somewhere??

OpenStudy (unklerhaukus):

if i am i cant see it ? where are you looking?

OpenStudy (anonymous):

\[\large R(t,y) = t^m y^n\] Right ??

OpenStudy (unklerhaukus):

oh, bother

OpenStudy (anonymous):

\[\large R(t,y) = t^{-2} y^{-3}\]

OpenStudy (unklerhaukus):

got my alphabet backwords

OpenStudy (unklerhaukus):

thanks

OpenStudy (anonymous):

Now you have to find f(t,y) again.. And I find f(t,y) for second one..

OpenStudy (anonymous):

No no you have found them correct I guess..

OpenStudy (unklerhaukus):

have i got them wrong/?

OpenStudy (anonymous):

\[f(t,y)=\int\limits R(t,y)N\text dy \implies f(t,y) = \int\limits(t^{-2}y^{-3})(-8t^3)dy\]

OpenStudy (anonymous):

No.. In finding f(t,y) you have done right..

OpenStudy (anonymous):

\[\implies -8t \int\limits (y^{-3})dy \implies f(t,y) = \frac{4t}{y^2} + g(t)\]

OpenStudy (anonymous):

May be but I tried..

OpenStudy (unklerhaukus):

\[f(t,y)=\int R(t,y)M\text dt\]\[=\int t^{-2}y^{-3}(2t^3y-y^3)\text dt\]\[=\int \left(2ty^{-2}-t^{-2}\right)\text dt\]\[=\frac{-2t^2}{y^{2}}+\frac{1}{2t}+g(y)\] \[f(t,y)=\int R(t,y)N\text dy\]\[=\int (t^{-2}y^{-3})(-2t^4) \text dy\]\[={-2}t^2\int y^{-3}\text dy\]\[=\frac{{-2}t^2}{y^{2}}+h(t)\] \[g(y)=0;\qquad h(t)=\frac 1{2t}\] \[f(t,y)=\frac{-2t^2}{y^{2}}+\frac{1}{2t}=c\]

OpenStudy (anonymous):

Oh yeah I think integral has to taken first for finding N..

OpenStudy (unklerhaukus):

N is −2t^4 from the original equation

OpenStudy (anonymous):

Yes..

OpenStudy (unklerhaukus):

we dont find N

OpenStudy (anonymous):

I have found N by integral of dN/dt sorry..

OpenStudy (anonymous):

\[=\int\limits (t^{-2}y^{-3})(-2t^4) \text dy \implies -2t^2 \int\limits y^{-3}dy = \frac{-2t^2}{(-2) \times y^2} = \frac{t^2}{y^2} + h(t)\] Don't you think so ??

OpenStudy (unklerhaukus):

ok so i have made some mistakes

OpenStudy (anonymous):

Mistakes are part of our lives.. Do it but try to do it once.. And after all we are now learners only.. Sometimes I cannot find mistakes of mine so Friends are always there to find your mistakes.. So, be friends forever so that you and I can find each other's mistake. Ha ha ha ha..

OpenStudy (unklerhaukus):

its a deal

OpenStudy (anonymous):

Okay as you said...

OpenStudy (unklerhaukus):

Ok my full solution; \[(2t^3y-y^3)\text dt-2t^4\text dy=0\] \[\frac{\partial M}{\partial y}=2t^3-3y^2;\qquad\qquad\frac{\partial N}{\partial t}=-8t^3\] \[\frac{\partial M}{\partial y}\neq\frac{\partial N}{\partial t}\] \[R=R(t,y)=t^my^n\] \[\frac{\partial MR(t,y)}{\partial y}=\frac{\partial NR(t,y)}{\partial t}\]\[\frac{\partial }{\partial y}\left((2t^3y-y^3)t^my^n\right)=\frac{\partial }{\partial t}\left((-2t^4)t^my^n\right)\]\[\frac{\partial }{\partial y}\left(2t^{m+3}y^{n+1}-t^my^{n+3}\right)=\frac{\partial }{\partial t}\left(-2t^{m+4}y^n\right)\] \[2(n+1)t^{m+3}y^{n}-(n+3)t^my^{n+2}=-2(m+4)t^{m+3}y^n\] \[2(n+1)=-2(m+4);\qquad\qquad (n+3)=0\]\[n+1=-m-4;\qquad\qquad n=-3\]\[-2=-m-4;\qquad\qquad m=-2\] \[R=R(t,y)=\frac1{t^{2}y^{3}}\] \[f(t,y)=\int R(t,y)M\text dt\]\[=\int \frac1{t^{2}y^{3}}(2t^3y-y^3)\text dt\]\[=\int \left(\frac{2t}{y^{2}}-\frac1{t^2}\right)\text dt\]\[=\frac{t^2}{y^{2}}+\frac{1}{t}+g(y)\] \[f(t,y)=\int R(t,y)N\text dy\]\[=\int \frac1{t^{2}y^{3}}(-2t^4) \text dy\]\[={-2}t^2\int \frac 1{y^3}\text dy\]\[=\frac{t^2}{y^{2}}+h(t)\] \[g(y)=0;\qquad\qquad h(t)=\frac 1t\] \[f(t,y)=\frac{t^2}{y^{2}}+\frac{1}{t}=c\]

OpenStudy (anonymous):

It is looking good to me.. How you put g(y) = 0 and h(t) = 1/t any logic or procedure ??

OpenStudy (unklerhaukus):

i found \(g,h\) because \[f(t,y)=\int R(t,y)M\text dt=\int R(t,y)N\text dy\]\[\qquad\qquad\frac{t^2}{y^{2}}+\frac{1}{t}+g(y)=\frac{t^2}{y^{2}}+h(t)\] comparing these

OpenStudy (anonymous):

Oh I see..

OpenStudy (unklerhaukus):

the other method is to take the partial of derivative of say \[\frac{t^2}{y^{2}}+\frac{1}{t}+g(y)\] and to hav it equal N

OpenStudy (anonymous):

I got this.. But now the solution remains or not ??

OpenStudy (unklerhaukus):

ah looks like the partial derivative should be equal to the integrating factor times N \[\frac{\partial }{\partial y}\left(\frac{t^2}{y^{2}}+\frac{1}{t}+g(y)\right)=\frac{-2t^4}{t^{2}y^{3}}\] \[-2\frac{t^2}{y^3}+g'(y)=\frac{-2t^2}{y^{3}}\] \[g'(y)=0\]

OpenStudy (unklerhaukus):

\[g(y)=\int 0\cdot \text dy=c\]

OpenStudy (anonymous):

yes..

OpenStudy (unklerhaukus):

\[\frac{\partial}{\partial t}\left(\frac{t^2}{y^{2}}+h(t)\right)=RM\]\[\frac{2t}{y^2}+h'(t)=\frac{2t^3y-y^3)}{t^{2}y^{3}}\]\[\frac{2t}{y^2}+h'(t)=\frac{2t}{y^2}-\frac 1{t^2}\]\[h'(t)=-\frac 1{t^2}\]\[h(t)=-\int \frac 1t\text dt\]\[h(t)=\frac 1t\]

OpenStudy (unklerhaukus):

\[\Large f(t,y)=\frac{t^2}{y^{2}}+\frac{1}{t}=c\]